OpenStudy (anonymous):

what is the quadratic function that is created with roots at 2 and 4 and a vertex at (3 1)

OpenStudy (tkhunny):

Well, we have the vertex form. y - k = a(x-h)^2 We are given the vertex (3,1) y - 1 = a(x-3)^2 We may wish to observe that 2-3 = -1 and 4-3 = 1. Thus, we have been given symmetric points, (2,0) and (4,0) and either should finish up for us. Substitute one of those points and solve for 'a'!

OpenStudy (anonymous):

im still so confused on this one, i may be overthinking it

OpenStudy (tkhunny):

Substitute one of those points into the equation with 'a' remaining. Show me what you get,

OpenStudy (anonymous):

y-1=2(x-3)^2?